# Characteristic initial value problems for the Einstein-Maxwell-scalar field equations in spherical symmetry

### v2

@@ III The characterisitic initial-boundary value problem
The formulation of the an initial boundary value problem for (II) - (20) is a consequence of the twice contracted Bianchi identities, $\nabla^{a}(E_{ab}-\frac{1}{2}g_{ab}(g^{cd}E_{cd}))$ and its electromagnetic counterpart $\nabla^{a}E_{a}$ Tamburino and Winicour (1966) and it follows the idea of Bondi et al. (1962); Sachs (1962a, b). The spirit is to pick from $(E_{ab},E_{a},E)$ a set of main equations to hold everywhere, then the twice contracted Bianchi identities and this set of main equation imply that the complement equations of the main equations either hold trivially as a consequence of the main equations or they must be satisfied at a given value of $\lambda=\lambda_{\mathcal{B}}=const$. The former is called trivial equation and the latter are the so-called supplementary equations. For the set (II) - (20), the following assignations are made:
@@
    which given by (16), (17), (19) and (20), respectively.
@@
    $E_{w\lambda}$ (27)
    represented by (LABEL:eq:E01).
The main equations (25) do not form a typical hierarchy for the field equations like in other characteristic formulations, see e.g. Bondi et al. (1962); Newman and Penrose (1962); Mädler and Winicour (2016). This is due to the appearence of the $r_{,w}$ derivative in (17).
Definition of the following new fields
@@
$\mathcal{L}$ $=$ $\frac{2r\left(r_{,\lambda}\Phi_{,w}-\rho\Phi_{,\lambda}\right)+\epsilon(Z+\lambda)\Phi_{,\lambda}}{r_{,\lambda}}$ (28c)
cast (25) into
@@
where we introduced the auxillary variable
@@
$\Phi(0,\lambda)=\mathcal{F}_{\Phi}(\lambda)$ (30)
on a null hypersurface $w=0$ for values $\lambda\geq\lambda_{\mathcal{B}}$, where $\mathcal{F}_{\Phi}(\lambda)$ is an arbitrary function. It further requires the following boundary conditions at $\lambda=\lambda_{\mathcal{B}}$ for all values $w>0$
@@
In particular, the boundary conditions for the variables $r$, $Q$, $\alpha$ and $V$ follow from physical considerations, while those for auxillary fields $Z$ and $\mathcal{L}$ they follow from the definitions (28) and are given by
@@
It seems “unpleasant” that the boundary conditions (32) for $Z$ and $\mathcal{L}$, contain $w$ derivatives of the data $\Phi$ as well as the conformal factor $r$ at the boundary $\lambda=\lambda_{\mathcal{B}}$. We will see in the next sections that this can be (partially) resolved by specifying the nature of the boundary hypersurface at $\lambda_{\mathcal{B}}$.
Since we can always translate the origin of an affine parameter by a transformation like $\lambda\rightarrow a(w)\lambda+b(w)$, we set hereafter $\lambda_{\mathcal{B}}=0$. In the following three sections we discuss the asymptotic initial value problem, the initial value problem at a world line and the double null initital boundary value problem for two intersecting null hypersurfaces.
@@ III.1 Asymptotic initial value problem for large values of the affine parameter
Let the scalar field have the following asymptotic expansion
@@
where the functions with subscipt indices in brackets labelling the order of $\lambda^{-i}$ are real functions depending on $w$. The function $\mathcal{L}$ is also decomposed into a real and imaginary part
$\mathcal{L}=L+iP$ (34)
@@
$\mathcal{J}=\frac{2(\varphi_{[1]}\psi_{[2]}-\varphi_{[2]}\psi_{[1]})}{\lambda^{4}}+O(\lambda^{-5})$ (35)
@@
where $H$, $R_{\infty}$, $Q_{\infty}$, $\alpha_{\infty}$, $Z_{\infty}$, $L_{\infty}$, $P_{\infty}$, $U_{\infty}$ and $V_{\infty}$ are functions of integrations depending on $w$. Using (17), we relate the functions integration $U_{\infty}$ and $V_{\infty}$
@@
$V_{\infty}$ $=$ $\frac{1-2\epsilon HR_{\infty}^{\prime}}{H^{2}}\;\;.$ (38)
The evolution equations (29g) relate the constants $L_{\infty}$ and $P_{\infty}$ to the $w$ derivatives of $\varphi_{[1]}$ and $\psi_{[1]}$, respectively,
@@
From this result, alternative calculation of the Bondi mass using the hypersurface variables is
@@
$x(w,\lambda)=H(w)\lambda+R_{\infty}(w)+O(\lambda^{-1})$ (44)
@@
which is the metric of an inertial observer for large values of the affine parameter. The coordinate frame in which $H=1$ is called Bondi frame and $w_{B}$ referred to as Bondi time corresponding to the inertial time of this observer. Consequently, in a regime where $H\neq 0$, it measures the redshift between asymptotic inertial observers and observes at finite affine parameter distance. In particular, if $H\neq 0$, we can always find a coordinate transformation of the type (46) such that $H=1$. If, however, $H=0$ for some time $w=w_{E}$, then $w_{E}$ corresponds to the time of formation of an event horizon, since the surface forming rays of the surface $dw=const$ do not have an endpoint in the asymptotic region. The redshift can also conveniently be calculated via
$H=\lim_{\lambda\rightarrow\infty}r_{,\lambda}$ (48)
@@
$\begin{split}r&=\lambda+R_{\infty}(w_{B})+O(\lambda^{-1})\\ V&=1-2\epsilon R_{\infty}^{\prime}(w_{B})+O(\lambda^{-2})\\ \end{split}$ (49)
and the scalar field $\Phi$ has the expansion,
$\Phi=\frac{\mathcal{C}(w_{B})}{r}+\frac{\mathcal{C}_{NP}}{r^{2}}+O(r^{-3})$ (50)
where $\mathcal{C}(w_{B})=\mathcal{C}_{\varphi}(w_{B})+i\mathcal{C}_{\psi}(w_{B})$ is the scalar monopole and $\mathcal{C}_{NP}$ is the conserved Newman-Penrose constant. The two can alternatively be determined via the limits
$\mathcal{C}$ $=$ $\lim_{r\rightarrow\infty}r\Phi=\lim_{\lambda\rightarrow\infty}H\lambda\Phi=H(\varphi_{[1]}+i\psi_{[1]})$ (51)
@@
$=$ $R_{\infty}\mathcal{C}+H^{2}(\varphi_{[2]}+i\psi_{[2]})$
@@
$\mathscr{N}(w_{B}):=\frac{d\mathcal{C}}{dw_{B}}\;\;$ (53)
is the news function of the scalar field, whose naming will become justified below. The relation between the Bondi time $w_{B}$, and the asymptotic time $w$ is given by asymptotic transformation (46) therefore
$\mathcal{N}(w_{B}):=\frac{d\mathcal{C}}{dw}\frac{dw}{dw_{B}}=H\frac{d\mathcal{C}}{dw}$ (54)
@@
$\begin{split}\lim_{\lambda\rightarrow\infty}\lambda^{3}\Psi_{2}=&\frac{2HZ_{\infty}-R_{\infty}}{2H^{3}}\\ &+\frac{\kappa\epsilon}{6H^{2}}\left[H\varphi_{[1]}\frac{d}{dw}H\varphi_{[1]}+H\psi_{[1]}\frac{d}{dw}H\psi_{[1]}\right]\end{split}$ (55)
in a Bondi frame ($H=1$, $r=\lambda$), we can deduce that
@@
$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left\{\left(\frac{d\varphi_{[1]}}{dw_{B}}\right)^{2}+\left(\frac{d\psi_{[1]}}{dw_{B}}\right)^{2}+2q\alpha_{\infty}\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]+q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)\right\}$ (58a)
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)+4\pi q\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]$ (58b)
which we can also write a in terms of the complex news function and complex scalar monopole
$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left|\mathcal{N}+iq\alpha_{\infty}\mathcal{C}\right|^{2}$ (59a)
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}|\mathcal{C}|^{2}-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)$ (59b)
As mentioned Sec. II, there is a gauge transformation to set [^3] Since $\alpha_{\infty}=\alpha_{\infty}(w_{B})$ this gauge transformation needs to be done on each null hypersurface $w_{B}=const$ for every value of $w_{B}$. $\alpha_{\infty}=0$. Using $\alpha_{\infty}=0$ in (59) casts the the Bondi mass conservation law and asymptotic the charge conservation law into the simple expressions
$\frac{dm_{B}}{dw_{B}}=\frac{\epsilon\kappa}{2}|\mathcal{N}|^{2}\;\;,\;\;\frac{dQ_{\infty}}{dw_{B}}=-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)\;\;$ (60)
From the definition of the news function $\mathcal{N}$ in (53) as the derivative of the scalar monopole $\mathcal{C}$ and (60), we deduce that the asymptotic initial value problem for the charged version ($q\neq 0$ ) requires specification of three scalar values and one arbitrary function depending on $w_{B}$ while the uncharged version $(q=0)$ only requires specification of two scalars and one function depending on $w_{B}$. In both of these cases, the free function is the news function $\mathcal{N}$. Its name arises from the same motivation as in the gravitational case, see (Bondi et al., 1962, Sec. 3), the function $\mathcal{N}$ determines the complete (temporal) behaviour of all the other physically measurable fields, $\mathcal{C}(w_{B})$, $m_{B}(w_{B})$ and $Q_{\infty}(w_{B})$.
Both of the two versions of these asymptotic initial value problems can be discussed together. Given a scalar news function $\mathcal{N}$ for $w_{B}\geq 0$. The three scalars are specified at the initial time $w_{B}=0$. They are $\mathcal{C}(0)$, $m_{B}(0)$ and $Q_{\infty}(0)$, which are the initial scalar monopole, the initial Bondi mass and the initial value of the global charge, respectively. The behavior of these fields for $w_{B}>0$ is then determined by the following hierarchy of ordinary differential equations of the asymptotic initial value problem
@@
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-2i\pi q\left[\mathcal{N}(w_{B})\bar{\mathcal{C}}(w_{B})-\bar{\mathcal{N}}(w_{B})\mathcal{C}(w_{B})\right]\;\;.$ (61c)
In the uncharged version, where the gauge charge $q$ vanishes, the last equation of (61) implies $Q_{\infty}(w_{B})=0$ as the initial value of the global charge $Q_{\infty}(0)$ is obviously zero. Thus only two initial values and one free function is required for the asymptotic initial value problem in this case.

### v3

@@ III The characterisitic initial-boundary value problem
The formulation of the CIBVP for (II) - (20) is a consequence of the twice contracted Bianchi identities, $\nabla^{a}(E_{ab}-\frac{1}{2}g_{ab}(g^{cd}E_{cd}))$ and its electromagnetic counterpart $\nabla^{a}E_{a}$ and it follows the idea of Bondi et al. (1962); Sachs (1962a, b); Tamburino and Winicour (1966). The spirit444 See e.g. Sachs (1962b); Tamburino and Winicour (1966) or Appendix A of Mädler (2019) for a discussion on the twice contracted Bianchi identity for a general metric adopted to a family of null hypersurfaces and the classification of main equations, supplementary equations and trivial equation. is to pick from $(E_{ab},E_{a},E)$ a set of main equations to hold everywhere, and then the twice contracted Bianchi identities and this set of main equations imply that the complementary equations of the main equations either hold trivially as a consequence of the main equations or they must be satisfied at a given value of $\lambda=\lambda_{\mathcal{B}}$=const. The former is called the trivial equation and the latter are the so-called supplementary equations. For the set (II) - (20), the following assignations are made:
@@
    which are given by (16), (17), (19), and (20), respectively.
@@
    $E_{w\lambda},$ (27)
    represented by (15).
In the Bondi-Sachs formulation of the CIBVP Bondi et al. (1962); Sachs (1962a); Mädler and Winicour (2016) the main equations form a natural hierarchical system of hypersurface equations that can be solved provided there is an initial profile of the dynamically propagating field(s). That is, each of the equations is solved one after another, in successive order. For the Einstein-scalar field equations in spherical symmetry while equipped with Bondi-Sachs coordinates, the only field propagating is the scalar field. There are two remaining fields in the system parametrizing the metric, the fields $\beta_{BS}$ and $V_{BS}$ (see, e.g., Gomez et al. (1994); Mädler and Gallo (2025)). Indeed, in that case [where $E_{\lambda}$ in (25) does not need to be considered], $E_{\lambda\lambda}$ is integrated first with the initial data to find $\beta_{BS}$. Then, with the initial data and $\beta_{BS}$, the field $V_{BS}$ is found by the integration of $q^{AB}E_{AB}$ in (25). Next, the wave equation $E$ in (25) is used to update the initial data from the initial null hypersurface to a null hypersurface at later times.
However, this beautiful and beneficial hierarchy for numerical integration is destroyed in the affine-null formulation, due to the appearance of the $r_{,w}$ derivative in (17). This additional time derivative is a consequence of the affine-null coordinate choice. The affine parameter $\lambda$ is related to the Bondi-Sachs area distance coordinate $r$ via $\lambda_{,r}(w,r,x^{A})=\exp[2\beta_{BS}(w,r,x^{A})]$. The transformation appears simple, but it actually comes along with a change in the $\partial_{w}$ directions555 The $\partial_{A}$-directions change in addition, if the assumption of spherical symmetry is dropped.. This results in the additional $w$-derivative in (17) Winicour (2013).
Nevertheless, we can establish the hierarchy in the affine-null case after the introduction of the following new fields
@@
$\mathcal{L}$ $=$ $\frac{2r\left(r_{,\lambda}\Phi_{,w}-\rho\Phi_{,\lambda}\right)+\epsilon(Z+\lambda)\Phi_{,\lambda}}{r_{,\lambda}}.$ (28c)
which cast (25) into
@@
where we introduced the auxiliary variable
@@
$\Phi(0,\lambda)=\mathcal{F}_{\Phi}(\lambda),$ (30)
on a null hypersurface $w=0$ for values $\lambda\geq\lambda_{\mathcal{B}}$, where $\mathcal{F}_{\Phi}(\lambda)$ is an arbitrary function. It further requires the following boundary conditions at $\lambda=\lambda_{\mathcal{B}}$ for all values $w\neq 0$
@@
In particular, the boundary conditions for the variables $r$, $Q$, $\alpha$, and $V$ follow from physical considerations, while those for auxiliary fields $Z$ and $\mathcal{L}$ follow from the definitions (28) and are given by
@@
The $r_{,w}$ derivative in the boundary condition for $Z$ in (32a) can be either removed or evaluated from $\Phi_{,w}$ via a differential equation in the formulation of the CIBVP (see next sections). Thus, the boundary conditions depend only on the $w$-derivative of the propagating physical field i.e., the derivative $\Phi_{,w}$. Consequently, the only propagating degree of freedom of the formulated CIBVP is the field that is subject to a hyperbolic equation, whose characteristics are given by the $w=$const null hypersurfaces. For the Einstein-Maxwell-scalar-field equations in spherical symmetry only the scalar field is determined by a hyperbolic partial differential equation, i.e., the wave equation $\mathcal{D}_{a}\mathcal{D}^{a}\Phi=0.$
Since we can always translate the origin of an affine parameter by a transformation like $\lambda\rightarrow a(w)\lambda+b(w)$, we set hereafter $\lambda_{\mathcal{B}}=0$. In the following three sections we discuss the asymptotic initial value problem, the initial value problem at a world line, and the double-null initial boundary value problem for two intersecting null hypersurfaces.
@@ III.1 Asymptotic initial value problem for large values of the affine parameter
To consider the asymptotic characteristic initial value problem, we assume that the fields have sufficient differentiability for large values of the affine parameter and in the limit $\lambda\rightarrow\infty$. An illustration of the null hypersurfaces for the parameter $\epsilon$ is given in Fig. 1. The scalar field is assumed to obey the radiation condition that the limit $\lim_{\lambda\rightarrow\infty}\lambda\Phi$ is finite on $w=$const null hypersurfaces.
Figure 1: The asymptotic null hypersurface for the cases $\epsilon=\pm 1$ are shown by spherical cuts at the location of “$\lambda=\infty$”. Either the idealized limit “$\lambda=\infty$” can be attached to the physical spacetime via the conformal Penrose compactification Penrose (1963) or it can be achieved by introducing a radial compactified coordinate, such as $\ell=1/\lambda$, and a proper regularization of the field equationMädler and Gallo (2025). The upper graphic illustrates a series of spherical cross sections of a Bondi frame, where the affine parameter $\lambda\rightarrow\infty$ on outgoing null hypersurfaces $w_{B}$=const. The Bondi time $w_{B}$ then corresponds to the retarded time $u_{B}$. The lower graphic shows the corresponding situation for ingoing null hypersurfaces toward a Bondi frame, where $\epsilon=+1$ and $w_{B}$ are ingoing null hypersurfaces. In this case, the Bondi time is the advanced coordinate $v_{B}$. In both graphics, coordinates $x^{A}$ parametrize the cuts at $w_{B}$=const and label the different rays (indicated by the arrow lines also carrying $\lambda\rightarrow\infty$ ) reaching the “$\lambda=\infty$” (null) hypersurface .
Then the scalar field has the following asymptotic expansion
@@
where the functions with subscript indices in brackets labeling the order of $\lambda^{-i}$ are real functions depending on $w$. The function $\mathcal{L}$ is also decomposed into real and imaginary parts
$\mathcal{L}=L+iP,$ (34)
@@
$\mathcal{J}=\frac{2(\varphi_{[1]}\psi_{[2]}-\varphi_{[2]}\psi_{[1]})}{\lambda^{4}}+O(\lambda^{-5}),$ (35)
@@
where $H$, $R_{\infty}$, $Q_{\infty}$, $\alpha_{\infty}$, $Z_{\infty}$, $L_{\infty}$, $P_{\infty}$, $U_{\infty}$ and $V_{\infty}$ are functions of integration depending on $w$. Using (17), we relate the functions integration $U_{\infty}$ and $V_{\infty}$,
@@
$V_{\infty}$ $=$ $\frac{1-2\epsilon HR_{\infty}^{\prime}}{H^{2}}\;\;,$ (38)
where denotes the total derivative with respect to the single independent argument of the respective function. The evolution equations (29g) relate the constants $L_{\infty}$ and $P_{\infty}$ to the $w$ derivatives of $\varphi_{[1]}$ and $\psi_{[1]}$, respectively,
@@
From this result, an alternative calculation of the Bondi mass using the hypersurface variables is
@@
$x(w,\lambda)=H(w)\lambda+R_{\infty}(w)+O(\lambda^{-1}),$ (44)
@@
which is the metric of an inertial observer for large values of the affine parameter. The coordinate frame in which $H=1$ is called the Bondi frame and $w_{B}$ is referred to as the Bondi time, corresponding to the inertial time of this observer. Consequently, in a regime where $H\neq 0$, it measures the redshift between asymptotic inertial observers and observers at finite affine parameter distance. In particular, if $H\neq 0$, we can always find a coordinate transformation of the type (46) such that $H=1$. If, however, $H=0$ for some time $w=w_{E}$, then $w_{E}$ corresponds to the time of formation of an event horizon, since the surface forming rays of the hypersurface $w=$const do not have an end point in the asymptotic region. The redshift can also conveniently be calculated via
$H=\lim_{\lambda\rightarrow\infty}r_{,\lambda}.$ (48)
@@
$\begin{split}r&=\lambda+R_{\infty}(w_{B})+O(\lambda^{-1}),\\ V&=1-2\epsilon R_{\infty}^{\prime}(w_{B})+O(\lambda^{-2}),\\ \end{split}$ (49)
and the scalar field $\Phi$ admits the expansion,
$\Phi=\frac{\mathcal{C}(w_{B})}{r}+\frac{\mathcal{C}_{NP}}{r^{2}}+O(r^{-3}),$ (50)
where $\mathcal{C}(w_{B})=\mathcal{C}_{\varphi}(w_{B})+i\mathcal{C}_{\psi}(w_{B})$ is the scalar monopole and $\mathcal{C}_{NP}$ is the conserved Newman-Penrose constant. These two quantities can alternatively be determined via the limits
$\mathcal{C}$ $=$ $\lim_{r\rightarrow\infty}r\Phi=\lim_{\lambda\rightarrow\infty}H\lambda\Phi$ (51)
$=$ $H(\varphi_{[1]}+i\psi_{[1]}),$
@@
$=$ $R_{\infty}\mathcal{C}+H^{2}(\varphi_{[2]}+i\psi_{[2]}).$
@@
$\mathcal{N}(w_{B}):=\frac{d\mathcal{C}}{dw_{B}},\;\;$ (53)
is the news function of the scalar field, whose naming will become justified below. The relation between the Bondi time $w_{B}$, and the asymptotic time $w$ is given by asymptotic transformation (46); therefore,
$\mathcal{N}(w_{B}):=\frac{d\mathcal{C}}{dw}\frac{dw}{dw_{B}}=H\frac{d\mathcal{C}}{dw}.$ (54)
@@
$\begin{split}\lim_{\lambda\rightarrow\infty}\lambda^{3}\Psi_{2}=&\frac{2HZ_{\infty}-R_{\infty}}{2H^{3}}\\ &+\frac{\kappa\epsilon}{6H^{2}}\left[H\varphi_{[1]}\frac{d}{dw}H\varphi_{[1]}+H\psi_{[1]}\frac{d}{dw}H\psi_{[1]}\right],\end{split}$ (55)
in a Bondi frame ($H=1$, $r=\lambda$), and we can deduce that
@@
$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left\{\left(\frac{d\varphi_{[1]}}{dw_{B}}\right)^{2}+\left(\frac{d\psi_{[1]}}{dw_{B}}\right)^{2}+2q\alpha_{\infty}\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right]+q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)\right\},$ (58a)
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}\left(\phi_{[1]}^{2}+\psi_{[1]}^{2}\right)+4\pi q\left[\phi_{[1]}\left(\frac{d\psi_{[1]}}{dw_{B}}\right)-\psi_{[1]}\left(\frac{d\phi_{[1]}}{dw_{B}}\right)\right],$ (58b)
which we can write also in terms of the complex news function and complex scalar monopole
$\frac{dm_{B}}{dw_{B}}$ $=$ $\frac{\epsilon\kappa}{2}\left|\mathcal{N}+iq\alpha_{\infty}\mathcal{C}\right|^{2},$ (59a)
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-4\pi q^{2}\alpha_{\infty}^{2}|\mathcal{C}|^{2}-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right).$ (59b)
As mentioned in Sec. II, there is a gauge transformation to set [^6] Since $\alpha_{\infty}=\alpha_{\infty}(w_{B})$ this gauge transformation needs to be done on each null hypersurface $w_{B}=$const for every value of $w_{B}$. $\alpha_{\infty}=0$. Using $\alpha_{\infty}=0$ in (59) casts the Bondi mass conservation law and the asymptotic charge conservation law into the simple expressions
$\frac{dm_{B}}{dw_{B}}=\frac{\epsilon\kappa}{2}|\mathcal{N}|^{2}\;\;,\;\;\frac{dQ_{\infty}}{dw_{B}}=-2i\pi q\left(\mathcal{N}\bar{\mathcal{C}}-\bar{\mathcal{N}}\mathcal{C}\right)\;\;.$ (60)
From the definition of the news function $\mathcal{N}$ in (53) as the derivative of the scalar monopole $\mathcal{C}$ and (60), we deduce that the asymptotic initial value problem for the charged version ($q\neq 0$ ) requires specification of three scalar values and one arbitrary function depending on $w_{B}$ while the uncharged version $(q=0)$ only requires specification of two scalars and one function depending on $w_{B}$. In both cases, the free function is the news function $\mathcal{N}$. Its name arises from the same motivation as in the gravitational case ( see (Bondi et al., 1962, Sec. 3)), and the function $\mathcal{N}$ determines the complete (temporal) behavior of all the other physically measurable fields, $\mathcal{C}(w_{B})$, $m_{B}(w_{B})$ and $Q_{\infty}(w_{B})$.
Both of the two versions of these asymptotic initial value problems can be treated together. Given a scalar news function $\mathcal{N}$ for $w_{B}\neq 0$, the three scalars are specified at the initial time $w_{B}=0$. They are $\mathcal{C}(0)$, $m_{B}(0)$ and $Q_{\infty}(0)$, which are the initial scalar monopole, the initial Bondi mass and the initial value of the global charge, respectively. The behavior of these fields for $w_{B}\neq 0$ is then determined by the following hierarchy of ordinary differential equations of the asymptotic initial value problem
@@
$\frac{dQ_{\infty}}{dw_{B}}$ $=$ $-2i\pi q\left[\mathcal{N}(w_{B})\bar{\mathcal{C}}(w_{B})-\bar{\mathcal{N}}(w_{B})\mathcal{C}(w_{B})\right]\;\;.$
In the uncharged version, where the gauge charge $q$ vanishes, the last equation of (61) implies $Q_{\infty}(w_{B})=0$ as the initial value of the global charge $Q_{\infty}(0)$ is obviously zero. Thus only two initial values and one free function are required for the asymptotic initial value problem in this case.